OpenStudy (anonymous):

Which point is a solution of y = 3x and x + 2y = –21? Select one: a. (–3, –9) b. (–9, –3) c. (3, –9) d. (–3, 9)

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

Use the x-values in the choices and plug them into the first equation, y = 3x/

OpenStudy (igreen):

Just plug in 3x for 'y' in the 2nd equation: x + 2y = -21 x + 2(3x) = -21 Can you simplify that? @boricua5

OpenStudy (anonymous):

So first, let's multiply -3 to it. -3 x 3 = ?

OpenStudy (anonymous):

@boricua5

OpenStudy (igreen):

x + 2(3x) = -21 x + 6x = -21 7x = -21 Now we divide 7 to both sides, what's -21 / 7? @boricua5

OpenStudy (anonymous):

@AnswerMyQuestions : \[-3 \times 3=9\]

OpenStudy (anonymous):

@iGreen : -21 / 7 = -3

OpenStudy (igreen):

Yes, so the x-value is -3, so the answer is either A or D. Now we can plug in x = -3 into any of the two equations. y = 3x y = 3(-3) What's 3 * -3?

OpenStudy (igreen):

@boricua5

OpenStudy (anonymous):

-9

OpenStudy (igreen):

Yep, so your answer is (-3, -9).

OpenStudy (anonymous):

thanks .